OpenStudy (anonymous):

can someone help me with this one please 16x^4+54x^2-7

OpenStudy (irishboy123):

1/8, -7/2

OpenStudy (anonymous):

How did you get that?

OpenStudy (irishboy123):

you didn't actually specify what you wanted doing with the equations! if you want stationary points or intercepts or whatever, you should spell that out. the numbers I gave solve 16y^2+54y-7 = 0 --- you should notice that i subbed y for x^2 and solved for **y**. THUS if you are trying to find the roots of the original question based in x, the roots are the sqrts of these numbers, and they number 4 because you gave a quartic. however, it could be that you are integrating the original equation or finding a volume of revolution or an arc length. hope this helps.

Directrix (directrix):

Are you looking for the solutions to this equation: 16x^4+54x^2-7=0 @Crazy.c.246

Directrix (directrix):

If so, factor: 16x^4+54x^2-7=0 (2x^2 + 7) * (8x^2 - 1) = 0